A) If 8^x+1 = 2^2x+7, what is x?
Answer #1

Pages Solution o (@+1) 8 ...(logas bloga) = 2(22+7) taking log both side we get (2+1) log 8 =(22+7) log 2 (4+1) Logest (2+1)